A thin circular metal disc of radius 500.0 mm is set rotating about a central axis normal to its plane. Upon raising its temperature gradually, the radius increases to 507.5 mm. The percentage change in the rotational kinetic energy will be

Grade:12

1 Answers

Arun
5 years ago
Dear Student
 
Rotational K.E. = L^2 / 2I
 
where L is angular momentu and I is momeny of inertia.
 
delta K /K = – delta I/ I = – 0.03
 
hence percentage increase = 3%